Integration by Partial Fractions


1.) ∫(8t^3 + 13)dt/(t+2)(4t^2+1)


2.) ∫(5x^2 -3x+ 18)dx/x(9-x^2) from 1 to 2


1
Expert's answer
2021-05-11T13:53:44-0400

Question 1:


Given


"\displaystyle\int \frac{8t^3+13}{(4t^2+1)(t+2)} dt"

Since the degree of the numerator is not less than the degree of the denominator, we perform long polynomial division to yield:



"\frac{8t^3+13}{(t+2)(4t^2+1)} = 2 + \frac{-16t^2-2t+9}{(4t^2+1)(t+2)}"

Next, perform partial fraction decomposition on the fraction:



"\frac{-16t^2-2t+9}{(4t^2+1)(t+2)}"

Let

"\frac{-16t^2-2t+9}{(4t^2+1)(t+2)}= \frac{At+B}{4t^2+1} + \frac{C}{t+2}"

Upon simplifying,



"-16t^2-2t+9 = (At+B)(t+2) + C(4t^2+1)"

Expand the right-hand side


"-16t^2-2t+9 = t^2(A+4C)+t(2A+B)+(2B+C)"

On comparing the coefficients of "x^2" , "x" and the constants terms on both sides, we obtain the following system of equations:



"-16 = A+4C; \quad -2 = 2A+B; \quad 9=2B+C"

Solving the three equations simultaneously, we get "A = -4, B=6" and "C=-3"


Therefore,



"\frac{-16t^2-2t+9}{(4t^2+1)(t+2)}= 2 + \frac{6-4t}{4t^2+1} + \frac{-3}{t+2}"

And



"\displaystyle\int \frac{8t^3+13}{(4t^2+1)(t+2)} dt = \displaystyle\int \left(2 + \frac{6-4t}{4t^2+1} -\frac{3}{t+2} \right) dt"

where,



"\displaystyle\int 2\, dt = 2t+c_1 ; \displaystyle\int \left( -\frac{3}{t+2} \right) dt = -3\ln|t+2| + c_2"

And


"\displaystyle\int \left( \frac{6-4t}{4t^2+1} \right) dt = \displaystyle\int \frac{6}{4t^2+1} \, dt -\displaystyle\int\frac{4t}{4t^2+2} dt"


Let "u=2t \Rightarrow \frac{du}{2}=dt." Similarly, let "v=4t^2+1 \Rightarrow \frac{dv}{2}=4t\,dt"


Therefore,



"\displaystyle\int \frac{6}{4t^2+1} \, dt -\displaystyle\int\frac{4t}{4t^2+2} dt = \displaystyle\int \frac{6}{2(u^2+1)} \, du -\displaystyle\int\frac{1}{2}\frac{1}{v} \,dv"

"= 3\tan^{-1} (u) -\frac{1}{2} \ln|v| + c_3"

"= 3\tan^{-1} (2t) -\frac{1}{2} \ln|4t^2+1| + c_3"

Upon combining the different integral results, we obtain



"\displaystyle\int \frac{8t^3+13}{(4t^2+1)(t+2)} dt = 2t+3\tan^{-1}(2t) -\frac{1}{2}\ln|4t^2+1|-3\ln|t+2| + C"

...................................................................................................................................................................................


Question 2:


Given

"\displaystyle\int_1^2 \frac{5x^2-3x+18}{x(9-x^2)} \, dx"

Factor the denominator and simplify the integrand as follows:



"\frac{5x^2-3x+18}{x(9-x^2)} = \frac{-5x^2+3x-18}{x(9-x^2)} = \frac{-5x^2+3x-18}{x(x-3)(x+3)}"

By partial fraction decomposition, let



"\frac{-5x^2+3x-18}{x(x-3)(x+3)} = \frac{A}{x} + \frac{B}{x-3}+\frac{C}{x+3}"



Simplify by multiplying both sides by "x(x-3)(x+3)." This gives



"-5x^2+3x-18 = x^2 (A+B+C) + 3(B-C)x - 9A"

On comparing the coefficients of "x^2" , "x" and the constants terms on both sides, we obtain the following system of equations:



"-5 = A+B+C ; \quad 3=3(B-C) ;\quad -18=-9A"

Upon solving these three equations simultaneously, we get "A=2, B=-3" and "C=-4"


Therefore,

"\frac{-5x^2+3x-18}{x(x-3)(x+3)} = \frac{2}{x} + \frac{-3}{x-3}+\frac{-4}{x+3}"

And

"\displaystyle\int_1^2\frac{-5x^2+3x-18}{x(x-3)(x+3)} \, dx = \displaystyle\int_1^2 \left(\frac{2}{x} + \frac{-3}{x-3}+\frac{-4}{x+3}\right)\, dx"


"=\left[ 2\ln|x| -3\ln|x-3| - 4\ln|x+3| \right]^2_1"

"=\left[ 2\ln(2) -3\ln(1) - 4\ln(5) \right] - \left[ 2\ln(1) -3\ln(2) - 4\ln(4) \right]"

"= 5\ln(2) - 4\ln(5)+ 4\ln(4) = 2.5732"